What is the mass of an object with a specific heat capacity of 750J/(kg∘C) that needs 5625J to have its temperature raised by 15.0∘C?
1
Expert's answer
2018-10-08T15:15:09-0400
We can find the mass of the object from the formula:
Q=mc∆t,
here, Q=5625 J is the amount of heat that required to raise the temperature of the object by ∆t=15.0 ℃, m is the mass of the object, c=750 J⁄(kg℃) is the specific heat capacity of the object.
Then, we get:
m=Q/c∆t=(5625 J)/(750 J/(kg℃)∙15.0 ℃)=0.5 kg.